What is the smallest of three consecutive integers if the sum of the smaller two integers is equal to 177 less than four times the largest?
Answer by CubeyThePenguin(3113)

consecutive integers: (x-1), x, (x+1)

(x-1) + x = 4(x+1) - 177
2x - 1 = 4x - 173
172 = 2x
x = 86

smallest integer = x = 86